Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 59     Entry time: 28 Feb 2023, 23:24
Author: Tokushi Sato 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, 1-352 pulses, 0.8 mJ, stable delivery

Optical laser delivery

  • 800 nm, 1-01 pulses, 282 kHz, stable delivery
  • ns laser for illumination

Achievements/Observations

  • Alignment for NKB focusing setup were finished
  • Al scattering signal was observed on AGIPD, need to move CSS with IRU-SLIT-1 opening
  • Dammit currently running on the "AGIPD" PC, automated data processing is on but will only process the first 100k frames.

Issues

  • No major issue to report
ELOG V3.1.4-7c3fd00